OpenStudy (anonymous):

Which of the following functions grows the fastest as x goes to infinity? 3^x ln(x) e^4x x^10

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@dan815 @Data_LG2 @dtan5457 @HelpBlahBlahBlah @HELPMEPLZ!! @jagr2713 @ganeshie8 @kropot72 @perl @Preetha @PRAETORIAN.10 @rational @SolomonZelman @sleepyhead314 @sleepyjess

OpenStudy (solomonzelman):

from fastest growth, to slowest growth. 1) exponential 2) x^(to some power) 3) logarithmic

OpenStudy (anonymous):

exponential r: 3^x and e^4x polynomial(x to some power): x^10 logarithmic: ln(x)

OpenStudy (anonymous):

I just dont know whether its 3^x and e^4x

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

you would agree that this function is \(\large\color{black}{ \displaystyle e^{4x} }\) is bigger than \(\large\color{black}{ \displaystyle 3^{x}}\). \(\large\color{black}{ \displaystyle e^{4x}~\Rightarrow~ \left(e^{4}\right)^x }\) and, that is biger than \(\large\color{black}{ \displaystyle 3^x }\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

very good. you don't need to tag me every second

OpenStudy (solomonzelman):

1. \(\large\color{black}{ \displaystyle e^{4x} }\) 2. \(\large\color{black}{ \displaystyle 3^{x} }\)

OpenStudy (solomonzelman):

then you tell me 3. and 4.

OpenStudy (anonymous):

3. x^10 4. ln(x)

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

any questions ?

OpenStudy (anonymous):

so this is in order of which is the fastest, yes?

OpenStudy (solomonzelman):

yes, from fastest to slowest e^(4x) 3^x x^10 ln(x)